OpenStudy (anonymous):

PLEASE HELP A juggler tosses a ball from her hand at a height of 1 meter with an initial upward velocity of 10 m/s. After how many seconds will the ball reach its highest point? I think the equation would be h= 4.9 m^2 + 10 t + 1 (I think)

OpenStudy (mrnood):

The equation you have written gives the HEIGHT at any time and is correct. However - the question asks you about time. The maximum height is gained when the velocity is 0 (i.e. it has stopped going up and is about to fall) the equaion you need is v = u + at v=0 u = 10m/s a= -9.8 m/s^2 so you can work out t

OpenStudy (anonymous):

equation should be h=- 4.9*t^2 m^2 + 10 t + 1 (I think)

OpenStudy (anonymous):

in general y = 1/2*at^2 + vt + y0 here a = -g, since acceleration due to gravity points downwards y = 1/2 (-9.81)*t^2 + 10*t + 1 y = -4.9t^2 + 10*t + 1

OpenStudy (anonymous):

as mr nood pointed out, you can use the equation vf = vi + at where vf = 0 vi = 10, a = -9.8 0 = 10 + -9.8*t